OpenStudy (anonymous):

the width of rectangle is fixed at 30cm. what lengths will make the perimeter greater than 76cm? i got 18cm?

OpenStudy (anonymous):

length>8

OpenStudy (anonymous):

p=2(l+w) p=2(l+30)

OpenStudy (anonymous):

\[2(l+30)>76\]

OpenStudy (anonymous):

2(30)+2(L)>76 60+2(L)>76 2(L)>16 L>8 simple honey

OpenStudy (anonymous):

\[l+30>38\]

OpenStudy (anonymous):

i ended up with 18..

OpenStudy (anonymous):

check mine it shows the step with the rectangle formula

OpenStudy (anonymous):

maybe i sub. wrong

OpenStudy (anonymous):

yea i divided wrong. it is eight. i always make small mistakes right at the end.

OpenStudy (anonymous):

thanks yall